Suppose the demand curve for product Y is given by Q = 300 – 2P + 4I, where I is income measured in thousands of dollars, P is price of product Y in $, Q is quantity of product Y. The supply curve is Q = 3P + 50, If I = 25. a) What is the price elasticity of demand for product Y at the equilibrium? b) What is the income elasticity of demand for product Y?


1
Expert's answer
2021-03-02T17:49:36-0500
"300-2p+100=3p+50"

"5p=350"

"p=70"

"Q=260"

a)


"E_p=-2\times \frac{70}{260}=-0.538"

b)


"E_I=4\times \frac {25}{260}=0.385"